# Symmetry breaking effects in pion couplings to constituent quarks currents

### v1

@@ Abstract
Pseudoscalar and axial neutral and charged pion-constituent quark coupling constants are investigated with nondegenerate quark masses in different kinematical points, off shell and on shell pions and constituent quarks. For the neutral pion, mixing effects are introduced by means of the pion mixing to states $P_{0}$ and $P_{8}$, that give rise to the $\pi^{0}-\eta-\eta^{\prime}$ meson mixing, and mixing of quark currents via corresponding mixing interactions. The relative behavior of charged and neutral pion coupling constants to quarks may be nearly the same - in the framework of the constituent quark model - as the pion-nucleon coupling constants if mixings are introduced. A very small pion coupling to strange quark current is also obtained. The dependence of the positive and negative pion-constituent quark coupling constant on the non-degeneracy of quark masses, for emission and absorption processes, is identified.
@@ 1 Introduction
Pions, as the (quasi)Goldsone bosons of Strong Interactions, have small masses as compared to baryons. As a consequence, pion couplings to other hadrons, especially to nucleons, have importance for a large number of low and intermediary energy processes. In spite of their small masses, they are well described in the framework of the constituent quark model (CQM) [1] in which a large binding energy is required. In part of these models, dynamical chiral symmetry breaking (DChSB) plays a role of one of the mechanisms of mass generation and contributes for large quark (effective) masses, although other mechanisms may contribute [2]. In most versions of the CQM, mesons interact with constituent quarks belonging to a baryon. These couplings must lead to the measurable meson-baryon coupling constants that parameterize such interactions phenomenologically. Several of the corresponding coupling constants might follow well known relations usually based in chiral symmetry and its dynamical breakdown such as Goldberger Treiman (GT) and partially conserved axial current relations [3, 4]. With the improvement of theoretical models and approaches and the increase of experimental precision, it became important to understand further effects of isospin and flavor symmetry breakings (ISB and FSB).
@@
There are several possible pion couplings to nucleons (or constituent quarks) with different Lorentz structures among which the axial and the pseudoscalar couplings are usually the most relevant ones. The corresponding coupling constants, and form factors, parameterize processes such as pion emission or absorption, besides being responsible for the Yukawa meson exchange among other processes. Isospin symmetry breaking effects in the pseudoscalar pion coupling to the nucleon have been investigated recently within different approaches [5, 6] leading to very similar conclusions that these effects must be very small. This is consistent with the fact that isospin symmetry breaking in flavor SU(2) QCD is ruled by up and down quark mass difference that is very small [7, 8, 9]. With different definitions of coupling constants, usually denoted by $f_{\pi^{a}NN^{\prime}}$ for $a=0,\pm$ and $N,N^{\prime}=n,p$, similar discrepancies between values of charged and neutral pion coupling constants were found, such as [6]: $g_{\pi^{0}nn}\simeq 13.397$, $g_{\pi^{0}pp}\simeq 13.227$ and $g_{\pi^{\pm}np}\simeq 13.226$. The pseudoscalar charged pion coupling constant is slightly smaller than the neutral one. In this work, the axial and pseudoscalar pion-constituent quark coupling constants will be derived from a quark determinant and investigated by considering non-degenerate up and down quark masses. Although the problem of establishing relations between quark and hadron level quantities might be firstly envisaged by means of group theoretical framework, the whole picture becomes more intricate due to confinement and to the dynamical character of Bethe-Salpeter and Faddeev type equations for bound quarks. Although this whole project may depend on the confinement mechanism and related properties, some aspects of hadron are seemingly independent of it according to phenomenology in which the CQM finds support for many hadrons. After all, color is confined, but flavor is not. With this program, the validity of the CQM description should be better defined and refined by investigating its consequences such as fine tuning (ISB or FSB) effects in meson couplings.
@@
In several versions of the CQM, the pion axial coupling is usually considered to be $g_{A}=1$, including in large Nc EFT proposed by Weinberg [10, 11], being the axial charge directly related to the nucleon spin [12]. Recently, a quark determinant was analyzed in the presence of local meson fields and constituent (background) quark currents [11, 13, 14]. In a large quark mass expansion, both pseudoscalar and axial couplings appear, besides other leading scalar and vector couplings with the possibility of higher order (non-leading) effective couplings. In this approach, by starting with a leading term of the QCD effective quark action for quark-antiquark interaction, by means of a (non-perturbative) gluon exchange, quark currents are split into quantum quark currents and background quark currents, that are dressed by components of a non perturbative gluon propagator leading to constituent quark currents. Quantum quark currents lead to meson fields, and to the chiral condensate, by means of the auxiliary field method. Mesons structure and dynamics arise in flavor multiplets of quark-antiquark components similarly to previous developments [15, 16, 17, 18]. In this approach, pseudoscalar meson mixings due to FSB have been identified [19] being possible to add to other known mechanism for meson mixing by means of instanton-induced interactions [20]. In the present work, flavor mixing is investigated in two ways, by means of mixings for $\pi^{0}-\eta-\eta^{\prime}$ mesons and flavor mixing effective quark-antiquark interactions, $G_{ij}(\bar{\psi}\Gamma\lambda_{i}\psi)(\bar{\psi}\Gamma\lambda_{j}\psi)$ for $\Gamma=i\gamma_{5}$, derived recently by a similar method [21, 22].
@@
Therefore, in the present work, an investigation of ISB (and FSB) in constituent quark-pion interactions is proposed by means of the pseudoscalar and axial pion couplings to constituent quark currents previously analyzed in the degenerate masses limit [11, 13]. Although the relation of these quark level couplings to the corresponding nucleon (or baryon in general) is not fully unambiguous for the reasons made explicit above, these quark level calculations and results should lead to more fine and precise assessment of to what extent, why and how constituent quark model works. Along these lines, the absolute values of the resulting coupling constants are not exactly the most relevant or interesting point of this work, but rather the effects of the nondegenerate quark masses in each channel, i.e. neutral or charged pions couplings to up or down quark currents, either pseudoscalar or axial. This might contribute to settle a route for analyzing hadron observable world with more complete grounds in low energy QCD. This work is organized as follows. In the next section some schematic explanation of the quark determinant in the presence of local meson fields and background quark currents, surrounded by a gluon cloud, is presented. The pseudoscalar and axial coupling constants are derived with the corresponding match equations. In Section (3) the possible mixings are analyzed for both pseudoscalar meson states and quark currents. For this, some mixing interactions obtained from flavor symmetry breaking in quark polarization [21, 22] are also considered. In Section (4) numerical results for the charged and neutral pion pseudoscalar and axial coupling constants are presented for different kinematical points: on shell and off shell (or massless) pions and constituent quarks. Some strangeness degree of freedom is shown to provide a small coupling to the neutral pion due to the mixings. In the last Section there is a Summary with conclusions.
@@ 2 Pion couplings from the quark determinant
The general procedure for the derivation of the quark determinant in the presence of background quark currents, surrounded by a gluon cloud represented by components of a gluon propagator, and local auxiliary meson fields displayed in quark-antiquark flavor U(3) nonets was developed in Refs. [11, 13, 14]. By selecting the (chiral) scalar and pseudoscalar auxiliary fields and by performing an usual chiral rotation, the scalar field nonet is eliminated in favor of non-linear parametrization of the (quasi) Goldstone boson octet as fluctuations around the vacuum. It can be written in terms of: $U=e^{\frac{iP\cdot\lambda}{F}}$ and $U^{\dagger}=e^{\frac{-iP\cdot\lambda}{F}}$, where $\lambda_{i}$ are the flavor Gell-Mann matrices and the pseudoscalar meson fields $P_{i}$ have canonical normalization. The quark determinant can be written as:
@@
The diagrammatic interpretation of the corresponding leading terms of the meson-quark interaction sector are shown in Fig. (1) and these diagrams - with one pion external leg - hold for both the pseudoscalar and the axial couplings with different form factors/coupling constants. Note that the flavor structure of the positive pion absorption coupling is different from the negative pion absorption but it is the same as the negative pion emission. In the expansion of the determinant, there also appear the scalar and vector (two-)pion couplings [13] that are of the same order of magnitude of the pseudoscalar and axial ones, although they present a considerably more intricate isospin structure and they will not be considered in this work.
@@
where $Tr_{F,D}$ is the trace in flavor and Dirac indices, $\tilde{S}_{0}^{f}(k)=(k^{2}-M_{f}^{2}+i\epsilon)^{-1}$. Although only the states $i,j=1,2,3$ will be worked out, some other states are related by mixings, $i,j=0,8$. For the definitions above, in the case of nondegenerate quark mass $G_{f_{1}\neq f_{2}}=G_{f_{2}\neq f_{1}}$. The difference is, however, very small being proportional to the up and down quark mass differences in the isospin subgroup. Due to U(1) electromagnetic invariance and CP, one has $G_{ij}=G_{ji}$ and $G_{11}=G_{22}$. Some complementary definitions that will be useful to work out mixings are presented in the Appendix (A). The positive and negative pion form factors present exactly the same form factors since $G_{11}$ is given by Eq. (A.1). In fact, it involves the average of positive or negative pion absorption and emission processes.
@@ 3 Quark currents, meson states and mixings
By writing interactions (6) explicitly in terms of charge eigenstates, for the pseudoscalar coupling and corresponding pion form factors, it yields:
@@ 3.1 Mixings
Another mixing term can be taken into account. For instance, there might appear neutral pions from to the oscillations of $\eta,\eta^{\prime}$ due to meson mixings. There are two different possible ways meson mixing can show up. First in processes that lead to pions, etas and etaprimes for which contamination from $\eta,\eta^{\prime}$ oscillations can lead to contributions for the neutral pion coupling to constituent quarks/nucleon. Second, in processes in which the original neutral pion oscillates to $\eta$ or $\eta^{\prime}$ that can couple to quarks/nucleons. Consider the following parametrization of these meson mixings (rotations):
@@
where the pion receives the smallest mixings and then $a_{3}<1$ with $a_{3}\sim 1$. These equations have been inspired in usual parametrization [26, 27]. Only two mixing terms will be needed below, $\eta-\pi^{0}$ or $\eta^{\prime}-\pi^{0}$ by means of $\bar{G}_{03}$ and $\bar{G}_{38}$. Furthermore, two types of interactions will be analyzed, the pseudoscalar and axial, so that all the coupling constants and mixing parameters will be labeled with an extra index $\xi=A,ps$ below. As a consequence, the following neutral pion coupling to quark currents, $J_{u}^{\xi},J_{d}^{\xi},J_{s}^{\xi}$, in both axial and pseudoscalar channels, are obtained:
@@
If these mixings were simply $\bar{G}_{03}^{\xi}=G_{03}^{\xi}$ and $\bar{G}_{38}^{\xi}=G_{38}^{\xi}$ the pion coupling to the strange-quark current would disappear, i.e. $G_{mix,ss}^{\xi}=0$.
@@ 3.2 Renormalization condition
To provide numerical results, the following renormalization condition will be considered by fixing the pseudoscalar charged pion coupling constant with a value based in phenomenology [28] (off shell pion and on shell quark):
@@
It turns out that, for the sets of parameters considered in this work, results will provide nearly $g_{A}(M^{2},0)^{\pi^{\pm}}\sim 1$ without further conditions. A different condition could be imposed by assuming each quark of the nucleon to be responsible for part of the couplings to mesons, such that $G_{ps}\sim 13/3$. However, for the condition above, pion-nucleon coupling constants should arise from an average of pion-constituent quark coupling constants. This might be considered in the Goldberber-Treiman (GT)-relation if the nucleon mass is considered instead of the constituent quark mass ($M_{q}$). Otherwise, with a constituent quark mass, GT-relation should be written as $\frac{G^{ps}}{g^{A}}=\frac{M_{q}}{f_{\pi}}(1+\Delta_{GT})$ being $\Delta_{GT}$ not small with respect to 1. Conversely, in other works [13, 30], an alternative condition was considered, i.e. to fix either $g_{A}=1$ or $g_{v}=12$ (vector meson coupling to constituent quarks) for the CQM.

### v2

# Symmetry breaking effects in pion couplings to constituent quark currents

@@ Abstract
Pseudoscalar and axial neutral and charged pion-constituent quark coupling constants are investigated with nondegenerate quark masses in different kinematical points, off shell and on shell pions and constituent quarks. By considering a large quark mass expansion of a quark determinant in the presence of local pion field and of constituent quark background currents, gluonic effects are considered by means of an effective gluon propagator that dresses quark currents. For the neutral pion, mixing effects are introduced by means of the pion mixing to states $P_{0}$ and $P_{8}$, that give rise to the $\pi^{0}-\eta-\eta^{\prime}$ meson mixing, and mixing of quark currents via corresponding mixing interactions. The relative behavior of charged and neutral pion coupling constants to quarks may be nearly the same - in the framework of the constituent quark model - as the pion-nucleon coupling constants if mixings are introduced. A very small pion coupling to strange quark current is also obtained. The dependence of the positive and negative pion-constituent quark coupling constant on the non-degeneracy of quark masses, for emission and absorption processes, is identified.
@@ 1 Introduction
Pions, as the (quasi)Goldsone bosons of Strong Interactions, have small masses as compared to baryons. As a consequence, pion couplings to other hadrons, especially to nucleons, have importance for a large number of low and intermediary energy processes. In spite of their small masses, they are well described in the framework of the constituent quark model (CQM) [1, 2] in which a large binding energy is required. In part of these models, dynamical chiral symmetry breaking (DChSB) plays a role of one of the mechanisms of mass generation and contributes for large quark (effective) masses, although other mechanisms may contribute [3]. In most versions of the CQM, mesons interact with constituent quarks belonging to a baryon. These couplings must lead to the measurable meson-baryon coupling constants that parameterize such interactions phenomenologically. Several of the corresponding coupling constants might follow well known relations usually based in chiral symmetry and its dynamical breakdown such as Goldberger Treiman relation (GTR) and partially conserved axial current [4, 5]. With the improvement of theoretical models and approaches and the increase of experimental precision, it became important to understand further effects of isospin and flavor symmetry breakings (ISB and FSB).
@@
There are several possible pion couplings to nucleons (or constituent quarks) with differerent Lorentz structures among which the axial and the pseudoscalar couplings are usually the most relevant ones. The corresponding coupling constants, and form factors, parameterize processes such as pion emission or absorption, besides being responsible for the Yukawa meson exchange among other processes. Isospin symmetry breaking effects in the pseudoscalar pion coupling to the nucleon have been investigated recently within different approaches [6, 7] leading to similar conclusions that these ISB effects must be very small. According to these works, the pseudoscalar neutral pion-neutron coupling constant is larger than the pseudoscalar neutral pion-proton coupling constant. However, it may be either slightly larger or considerably smaller than the charged pion coupling constant. Nevertheless, the resulting overall small ISB is consistent with the fact that isospin symmetry breaking in flavor SU(2) QCD is ruled by up and down quark mass difference that is very small [8, 9, 10]. Other related consequences have been exploited in the literature, for example in [6, 11].
@@
In several versions of the CQM, the pion axial coupling is usually considered to be $g_{A}=1$, including in large Nc EFT proposed by Weinberg [12, 13], being the axial charge directly related to the nucleon spin [14]. In the CQM, meson couplings to baryons, for example, are decomposed in meson couplings to constituent quarks [15]. Recently, a quark determinant was analyzed in the presence of local meson fields and constituent (background) quark currents [13, 16, 17]. In a large quark mass expansion, both pseudoscalar and axial couplings appear, besides other leading scalar and vector couplings with the possibility of higher order (non-leading) effective couplings. In this approach, by starting with a leading term of the QCD effective quark action for quark-antiquark interaction, by means of a (nonperturbative) gluon exchange, quark currents are split into quantum quark currents and background quark currents, that are dressed by components of a non perturbative gluon propagator leading to constituent quark currents. Quantum quark currents lead to meson fields, and to the chiral condensate, by means of the auxiliary field method. Mesons structure and dynamics arise in flavor multiplets of quark-antiquark components similarly to previous developments [18, 19, 20, 21]. In this approach, pseudoscalar meson mixings due to FSB have been identified [22] being possible to add to other known mechanism for meson mixing by means of instanton-induced interactions [23]. In the present work, flavor mixing are investigated in two ways, by means of mixings for $\pi^{0}-\eta-\eta^{\prime}$ mesons and flavor mixing at the level of quark flavor currents in effective quark-antiquark interactions, $G_{ij}(\bar{\psi}\Gamma\lambda_{i}\psi)(\bar{\psi}\Gamma\lambda_{j}\psi)$ for $\Gamma=i\gamma_{5}$, that have been derived by a similar method [24, 25].
In this work, the axial and pseudoscalar pion-constituent quark coupling constants will be derived from a quark determinant and investigated by considering nondegenerate up and down quark masses. The longstanding problem of establishing relations between quark and hadron level quantities are usually firstly envisaged by means of group theoretical framework although the whole picture might become more intrincated due to several involved issues such as confinement and to the dynamical character of Bethe-Salpeter and Faddeev type equations for bound quarks. Although this whole project may depend on the confinment mechanism and related properties, some aspects of hadron are seemingly independent of it according to phenomenology in which the CQM finds support for many hadrons. Afterall, color is confined but flavor is not. With this program, the validity of the CQM description should be better defined and refined by investigating its consequences such as fine tuning (ISB or FSB) effects in meson couplings.
@@
Therefore, in the present work, an investigation of the ISB (and FSB) in constituent quark-pion interactions is proposed by means of the pseudoscalar and axial pion couplings to constituent quark currents previously analyzed in the degenerate masses limit [13, 16]. Although the relation of these quark level couplings to the corresponding nucleon (or baryon in general) is not fully unambiguous for the reasons explicitated above, these quark level calculations and results should lead to more fine and precise assessment of to what extent, why and how constituent quark model works. Along these lines, the absolute values of the resulting coupling constants are not exactly the most relevant or interesting point of this work, but rather the effects of the nondegenerate quark masses in each channel, i.e. neutral or charged pions couplings to up or down quark currents, either pseudoscalar or axial. This might contribute to settle a route for analyzing hadron observable world with more complete grounds in low energy QCD. As pointed out in [16] momentum dependency of pseudoscalar form factor within this approach is not reasonable due to the momentum independent quark mass. Coupling constants, however, are well described. Since this level of calculation does not provide the running quark mass, the momentum dependency of form factors will not be investigated in this work. The present calculation has a dynamical character since it corresponds to deriving an effective action for pions, being mesons arranged in flavor (and spin) multiplets, and constituent quark currents by articulating the quark determinant obtained from a quark-antiquark interaction mediated by dressed gluons. The pion-constituent quark coupling constants will be calculated at different kinematical points from the corresponding dynamical form factors. This work is organized as follows. In the next section some schematic explanation of the quark determinant, and its large quark mass expansion, in the presence of local meson fields and background quark currents, surrounded by a gluon cloud, is presented. The pseudoscalar and axial coupling constants are derived with the corresponding match equations. In Section (3) the possible mixings are analyzed for both pseudoscalar meson states and quark currents. For this, some mixing interactions obtained from flavor symmetry breaking in quark polarization [24, 25] are also considered. In Section (4) numerical results for the charged and neutral pion pseudoscalar and axial coupling constants are presented for different kinematical points: on shell and off shell (or massless) pions and constituent quarks. Some strangeness degree of freedom is shown to provide a small coupling of the strangeness quark currents to the neutral pion due to the flavor mixings. In the last Section there is a Summary with conclusions.
@@ 2 Pion couplings from the quark determinant
The general procedure for the derivation of the quark derminant in the presence of background quark currents, surrounded by a gluon cloud represented by components of a gluon propagator, and local auxiliary meson fields displayed in quark-antiquark flavor U(3) nonets was developed in Refs. [13, 16, 17]. By selecting the (chiral) scalar and pseudoscalar auxiliary fields and by perfoming an usual chiral rotation, the scalar field nonet is eliminated in favor of non-linear parameterization of the (quasi) Goldstone boson octet as fluctuations around the vacuum. It can be written in terms of: $U=e^{\frac{iP\cdot\lambda}{F}}$ and $U^{\dagger}=e^{\frac{-iP\cdot\lambda}{F}}$, where $\lambda_{i}$ are the flavor Gell-Mann matrices and the pseudoscalar meson fields $P_{i}$ have canonical normalization. The quark determinant can be written as:
@@
The diagrammatic interpretation of the corresponding leading terms of the meson-quark interaction sector are shown in Fig. (1) and these diagrams - with one pion external leg - hold for both the pseudoscalar and the axial couplings with different form factors/coupling constants. Note that the flavor structure of the positive pion absorption coupling is different from the negative pion absorption but it is the same as the negative pion emission. In the expansion of the determinant, there also appear the scalar and vector (two-)pion couplings [16] that are of the same order of magnitude of the pseudoscalar and axial ones, although they present a considerably more intrincate isospin structure and they will not be considered in this work.
@@
where $Tr_{F,D}$ is the trace in flavor and Dirac indices, $\tilde{S}_{0}^{f}(k)=(k^{2}-M_{f}^{2}+i\epsilon)^{-1}$. Although only the states $i,j=1,2,3$ will be worked out, some other states are related by mixings, $i,j=0,8$. For the definitions above, in the case of nondegenerate quark mass $G_{f_{1}\neq f_{2}}=G_{f_{2}\neq f_{1}}$. The difference is, however, very small being proportional to the up and down quark mass differences in the isospin subgroup. Due to U(1) electromagnetic invariance and CP, one has $G_{ij}=G_{ji}$ and $G_{11}=G_{22}$. Some complementary definitions that will be useful to work out mixings are presented in the Appendix (A). Positive and negative pions present exactly the same form factors since $G_{11}$ and $G_{22}$ are given by Eq. (A.1). In fact, it involves the average of positive or negative pion absorption and emission processes. Note that the equations for the axial and pseudoscalar pion couplings are independent of each other in the approach considered in the present work, and none of them is calculated by considering the GTR which, nonetheless, is satisfied quite well when comparing numerical results. It is also interesting to emphasize that, as a consequence of the standard methods applied for the derivation of the quark determinant, pion couples directly to the pseudoscalar and axial quark currents. It is worth to emphasize that the above form factors were obtained in a dynamical way and not from the QCD vertex function as calculated diagrammatically or in a Schwinger-Dyson equations framework. Therefore coupling constants do not emerge as the residue of the pole from the corresponding Bethe-Salpeter equation (BSE) [29, 30]. The relation between the present approach and the standard one based in BSE has not yet been settled.
@@ 3 Quark currents, meson states and mixings
By writing interactions (6) explicitely in terms of charge eigenstates, for the pseudoscalar coupling and corresponding pion form factors, it yields:
@@ 3.1 Mixings
Another mixing term can be taken into account. For instance, there might appear neutral pions from to the oscillations of $\eta,\eta^{\prime}$ due to meson mixings. There are two different possible ways meson mixing can show up. First in processes that lead to pions, etas and etaprimes for which contaminations from $\eta,\eta^{\prime}$ oscillations can lead to contributions for the neutral pion coupling to constituent quarks/nucleon. Second, in processes in which the original neutral pion oscillates to $\eta$ or $\eta^{\prime}$ that can couple to quarks/nucleons. Consider the following parameterization of these meson mixings (rotations):
@@
where the pion receives the smallest mixings and then $a_{3}<1$ with $a_{3}\sim 1$. These equations have been inspired in usual parameterizations [31, 32]. Only two mixing terms will be needed below, $\eta-\pi^{0}$ or $\eta^{\prime}-\pi^{0}$ by means of $\bar{G}_{03}$ and $\bar{G}_{38}$. Furthermore, two types of interactions will be analyzed, the pseudoscalar and axial, so that all the coupling constants and mixing parameters will be labeled with an extra index $\xi=A,ps$ below. As a consequence, the following neutral pion coupling to quark currents, $J_{u}^{\xi},J_{d}^{\xi},J_{s}^{\xi}$, in both axial and pseudoscalar channels, are obtained:
@@
If these mixings were simply $\bar{G}_{03}^{\xi}=G_{03}^{\xi}$ and $\bar{G}_{38}^{\xi}=G_{38}^{\xi}$ the pion coupling to strange quark current would disappear, i.e. $G_{mix,ss}^{\xi}=0$.
@@ 3.2 Renormalization condition
To provide numerical results, the folllowing renormalization condition will be considered by fixing the pseudoscalar charged pion coupling constant with a value based in phenomenology [33] (off shell pion and on shell quark):
@@
It turns out that, for the sets of parameters considered in this work, results will provide nearly $g_{A}(M^{2},0)^{\pi^{\pm}}\sim 1$ without further conditions. Conversely, in other works [16, 34], an alternative condition was considered, i.e. to fix either $g_{A}=1$ or $g_{v}=12$ (vector meson coupling to constituent quarks) for the CQM. A different condition could be imposed by assuming each quark of the nucleon to be responsible for part of the couplings to mesons along other investigations [35], such that $G_{ps}\sim 13/3$. However, for the condition above, pion-nucleon coupling constants should arise from an average of pion-constituent quark coupling constants instead as a sum. This might be considered in the Goldberber-Treiman (GT)-relation if the nucleon mass is considered instead of the constituent quark mass ($M_{q}$). Otherwise, with a constituent quark mass, GT-relation should be written as $\frac{G^{ps}}{g^{A}}=\frac{M_{q}}{f_{\pi}}(1+\Delta_{GT})$ being $\Delta_{GT}$ not small with respect to 1.